# Smallness of neutrino masses and leptogenesis in

### v1

@@ 4 Conclusions
We argued that the observed mass hierarchy in the lepton sector can be caused by the approximate discrete $Z_{2}$ symmetry. Such $Z_{2}$ symmetry may also lead to three generations of relatively light composite neutral fermions and fermions with charges $\pm 1$. These resonances compose $SU(2)_{W}$ doublets $\tilde{L}_{2i}$ and can have masses within the $1-2\,\mbox{TeV}$ range so that they can be discovered at the LHC in the near future. The couplings of $\tilde{L}_{2i}$ to the SM leptons are quite suppressed because they are defined by the small parameters $s_{2i}\lesssim 10^{-5}$. It was shown that in this case the observed baryon asymmetry can be induced if the particle spectrum of the CHM3 contains an additional elementary Majorana fermion $n$ with mass around $10\,\mbox{TeV}$. The corresponding lepton asymmetry is generated due to the out–of equilibrium decays of $n$ into $\tilde{L}_{2i}$ and Higgs doublet even when CP is almost preserved.

### v2

@@ 4 Conclusions
We argued that the observed mass hierarchy in the lepton sector can be caused by the approximate $U(1)_{L}$ and discrete $Z_{2}$ symmetries. Such $Z_{2}$ symmetry may also lead to three generations of relatively light composite neutral fermions and fermions with charges $\pm 1$. These resonances compose $SU(2)_{W}$ doublets $\tilde{L}_{2i}$ and can have masses within the $1-2\,\mbox{TeV}$ range so that they can be discovered at the LHC in the near future. The couplings of $\tilde{L}_{2i}$ to the SM leptons are quite suppressed because they are defined by the small parameters $s_{2i}\lesssim 10^{-5}$. It was shown that in this case the observed baryon asymmetry can be induced if the particle spectrum of the CHM3 contains an additional elementary Majorana fermion $n$ with mass around $10\,\mbox{TeV}$. The corresponding lepton asymmetry may be generated due to the out–of equilibrium decays of $n$ into $\tilde{L}_{2i}$ and Higgs doublet even when CP is almost preserved.